OpenStudy (anonymous):

Solve for x. \[x^4( {18 \over x^4} + {1 \over x^2})=4(x^4)\]

OpenStudy (anonymous):

\[4x^4 - x^2 -18 =0\]

OpenStudy (anonymous):

this has real and complex roots

OpenStudy (ash2326):

@order put \(x^2=u\) then \(x^4=u^2\) Now our equation becomes a quadratic equation \[4u^2-u-18=0\] Can you solve now?

OpenStudy (anonymous):

Yes I think so. Let me try.

OpenStudy (anonymous):

Got it, thanks :)

OpenStudy (anonymous):

(18/x^2)+(1) = 4x^2 18 + x^2 = 4 x^4 4x^4 - x^2 - 18 = 0 (x^2 + 2)(4x^2 - 9) x^2 = -2 OR x^2 = 9/4 ==> x = +- 3/2